OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 5 to the fourth power, over 25 = 52 P.S. I'm extremely bad at math

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{5^4}{25}=5^2\] right ?

OpenStudy (misty1212):

\[\frac{5^4}{25}=\frac{4^4}{5^2}=5^{4-2}=5^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much